OpenStudy (anonymous):

How long does it take for the plane to stop with this acceleration? Answer in units of s

Parth (parthkohli):

\[v^2 = u^2 + 2as\]I see only one equation of motion with \(s\).

Parth (parthkohli):

Anyway, what is the acceleration?

OpenStudy (anonymous):

not sure if he means units of "seconds"

OpenStudy (anonymous):

africanboy you need to provide velocity and acceleration(values)

Parth (parthkohli):

The question is incomplete :|

OpenStudy (anonymous):

i will agree with that

OpenStudy (anonymous):

A plane landing on a small tropical island has just 64 m of runway on which to stop. If its initial speed is 66 m/s, what is the maximum acceleration of the plane during landing, assuming it to be constant? Answer in units of m/s How long does it take for the plane to stop with this acceleration? Answer in units of s

OpenStudy (anonymous):

Fundamental kinematics equations to solve for time: \[x=x_0+v_0t+\frac{1}{2}at^2, \space v=at\]

OpenStudy (anonymous):

The equation Parth provided is a combination of the above two and can be used to find the acceleration.

OpenStudy (anonymous):

i just need help with the second part

OpenStudy (anonymous):

do you already know the acceleration?

OpenStudy (anonymous):

Then use the first fundamental kinematics equation.

OpenStudy (anonymous):

It's a quadratic equation in t, and you can use the quadratic formula to solve.

OpenStudy (anonymous):

the acceleration is -34.03125

OpenStudy (anonymous):

Actually, I think the second one will work too...

OpenStudy (anonymous):

0=1/2(-34.03125)t^2+66t-64 solve for t

OpenStudy (anonymous):

yes, the second equation, v=at, works too and is a lot easier.

OpenStudy (anonymous):

thanks this is my first time using this website

OpenStudy (anonymous):

I got 64/33 s. How about you?

OpenStudy (anonymous):

Though, you'll probably want to round that to 2 significant figures to be honest about the uncertainty.

OpenStudy (anonymous):

sorry still cant find t

OpenStudy (anonymous):

\[\Delta v = a \cdot t\] Is the simplest way to get it. You have the acceleration, a, and the change in velocity is (0m/s - 66m/s) = -66m/s, so use that info to solve for t.